OpenStudy (anonymous):

3x*x-2x+16=0, solve for x. I don't understand what to do with all the x's!

OpenStudy (anonymous):

3x^2 - 2x + 16 = 0 (3x - 8)(x + 2) = 0 That's the first part

OpenStudy (anonymous):

thanks dude, that makes mor sense, i kept trying to divd by x